# AttentionStore: Cost-effective Attention Reuse across Multi-turn Conversations in Large Language Model Serving

### v1

@@ 4.3.4 Performance of Decoupled KV Cache Truncation
When the context window exceeds its limit, AttentionStore truncates the KV cache directly, thus avoiding the need for re-computation and reducing overhead. We evaluate the effectiveness of the way AS used to manage the context overflow. Specifically, we compare a baseline approach overflow (OF) that embeds positional encoding within the KV caches, leading to the invalidation of KV caches in the KV caching system. OF relegates context overflow to be managed by recomputation. This experiment uses 128GB DRAM and 10TB SSD. As evident from Figure 22(a), comparing OF with AS, the hit rates decrease by 52%, 56%, 57%, 48% for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. This decline is attributed to the fact that if applying OF, every instance of context overflow necessitates context truncation, thereby invalidating the KV caches in the KV caching system. This decrease in hit rate subsequently translates to the longer GPU time as shown in Figure 22(b).
@@ 4.3.6 Performance for Long Sequence Inference
4.3.7 The Cache Capacity Requirement
(a) Impact on hit rate.
(b) Impact on throughput.
4.3.8 Impact of Caching Storage Mediums
(a) Impact of caching storage mediums on hit rates.
(b) Impact of caching storage mediums on GPU time.
Figure 26: Performance under various caching configurations.
Some existing works [66, 19, 30] employ only the HBM space for caching the KV caches of multi-turn conversations. We here compare the performance of mechanism caching KVs on HBMs with that of AttentionStore caching KVs on DRAM and SSDs. In the experiments, we configure the size of the HBM cache as 10GB, the size of DRAM as 128GB, and the size of SSDs as 10TB. Figure 26(a) shows cache hit rates and inference performance of different mechanisms. The hit rate of the HBM-only caching method is nearly 0% for all models due to the limited capacity of HBM. Using HBM with DRAM improves the cache hit rate to 1.9%, 0.9%, 4.4%, and 11.2% for models LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, by further extending the cache capacity with SSDs, AttentionStore improves the cache hit rate to 76%, 56%, 87%, and 89% for models LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. With higher hit rates, AttentionStore significantly improves the inference performance compared to the HBM-only/HBM+DRAM policies as shown in Figure 26(b).
@@ 6 Conclusion
This paper proposes AttentionStore, a new attention that allows the reuse of the KV caches for any ensuing turns of the same conversation, achieving a significant reduction in the recomputation overhead of KV caches in LLMs. To improve the efficiency of AttentionStore, we design overlapped KV cache access, hierarchical KV cache placement, and positional encoding decoupled KV cache truncation schemes. Extensive experimental results demonstrate that AttentionStore significantly decreases the TTFT by up to 88% and improves the prompt prefilling throughput by 8.2$\times$ for multi-turn conversations. It reduces the end-to-end inference cost by up to 56%. It also decreases the TTFT by up to 95% and enhances prompt prefilling throughput by 22$\times$ for long sequence inference.

### v2

@@ 4.3.4 Performance of Decoupled KV Cache Truncation
When the context window exceeds its limit, AttentionStore truncates the KV cache directly, thus avoiding the need for re-computation and reducing overhead. We evaluate the effectiveness of the way AS used to manage the context overflow. Specifically, we compare a baseline approach overflow (OF) that embeds positional encoding within the KV caches, leading to the invalidation of KV caches in the KV caching system. OF relegates context overflow to be managed by recomputation. This experiment uses 128GB DRAM and 10TB SSD. As evident from Figure 22(a), comparing OF with AS, the hit rates decrease by 17.6%, 41.5%, 18.1%, 18.4% for LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. This decline is attributed to the fact that if applying OF, every instance of context overflow necessitates context truncation, thereby invalidating the KV caches in the KV caching system. This decrease in hit rate subsequently translates to the longer GPU time as shown in Figure 22(b).
@@
4.3.5 The Cache Capacity Requirement
(a) Impact on hit rate.
(b) Impact on throughput.
4.3.6 Impact of Caching Storage Mediums
(a) Impact of caching storage mediums on hit rates.
(b) Impact of caching storage mediums on GPU time.
Figure 24: Performance under various caching configurations.
Some existing works [67, 19, 30] employ only the HBM space for caching the KV caches of multi-turn conversations. We here compare the performance of mechanism caching KVs on HBMs with that of AttentionStore caching KVs on DRAM and SSDs. In the experiments, we configure the size of the HBM cache as 10GB, the size of DRAM as 128GB, and the size of SSDs as 10TB. Figure 24(a) shows cache hit rates and inference performance of different mechanisms. The hit rate of the HBM-only caching method is nearly 0% for all models due to the limited capacity of HBM. Using HBM with DRAM improves the cache hit rate to 3.4%, 1.7%, 7.7%, and 19.1% for models LLaMA-13B, LLaMA-65B, LLaMA-70B and Falcon-40B, respectively. In contrast, by further extending the cache capacity with SSDs, AttentionStore improves the cache hit rate to 86%, 71%, 89%, and 90% for models LLaMA-13B, LLaMA-65B, LLaMA-70B, and Falcon-40B, respectively. With higher hit rates, AttentionStore significantly improves the inference performance compared to the HBM-only/HBM+DRAM policies as shown in Figure 24(b).
@@ 6 Conclusion
This paper proposes AttentionStore, a new attention that allows the reuse of the KV caches for any ensuing turns of the same conversation, achieving a significant reduction in the recomputation overhead of KV caches in LLMs. To improve the efficiency of AttentionStore, we design overlapped KV cache access, hierarchical KV cache placement, and positional encoding decoupled KV cache truncation schemes. Extensive experimental results demonstrate that AttentionStore significantly decreases the TTFT by up to 87% and improves the prompt prefilling throughput by 7.8$\times$ for multi-turn conversations. It reduces the end-to-end inference cost by up to 70%. It also decreases the TTFT by up to 95% and enhances prompt prefilling throughput by 22$\times$ for long sequence inference.